# Configuration Interaction Guided Sampling with Interpretable Restricted Boltzmann Machine

### v1

@@ 2.2 Configuration Interaction
This exponential growth in the number of determinants with system size renders FCI intractable for large molecules or complex materials. While FCI provides the most accurate solution within a chosen basis set, exploring this vast configuration space becomes computationally intractable.
To overcome this prohibitive configuration space challenge, recent advancements in the field of machine learning offer promising avenues. Techniques like Restricted Boltzmann Machines (RBMs) can be employed as generative models [20, 21]. These models can learn the underlying probability distribution of the electronic system and efficiently sample the most relevant configurations for the FCI calculation. This selective approach, guided by the RBM, significantly reduces the computational cost compared to traditional methods that explore the configuration space randomly. The next section will delve into how RBMs can be harnessed to address the limitations of FCI and pave the way for studying electronic structures in complex systems.
2.3 Restricted Boltzmann Machine
@@
RBM (Restricted Boltzmann Machines) is a type of generative model and unsupervised learning algorithm used in the field of machine learning. RBM is a probabilistic energy-based neural network consisting of two layers of nodes: a visible layer and a hidden layer. The nodes in the visible layer represent observed variables, while the nodes in the hidden layer represent latent or unobserved features. These two layers are fully connected to each other, but there are no connections within each layer (See Figure 1).
Figure 1: General architecture of the Restricted Boltzmann Machines (RBM).
@@
The main goal of an RBM is to learn the synaptic weights that maximize the joint probability of the observed data in the visible layer and the latent representations in the hidden layer, or in other words the RBM as a generative model, represents a probability distribution (over states of the visible units) where low-energy configurations have higher probability. The energy is determined by the connection weights which are the parameters to be learnt from the data to built a generative model that learns the underlaying dynamic of the system [22]. This is achieved using an algorithm called “Gibbs sampling” or “Contrastive Divergence” [23].
@@
Because there are no direct connections between hidden units in an RBM, it is not necessary to know all the system configurations but rather the probabilities that connect a visible state to each of the nodes in the hidden state. Therefore, obtaining a sample of $\langle v_{i}h_{j}\rangle_{data}$ is very easy. Given a randomly selected training input, $v$, the binary state, $h_{j}$, of each hidden unit, $j$, is set to 1 with the following probability [24]:
@@
The problem arises when trying to obtain the expectation under the model distribution $\langle v_{i}h_{j}\rangle_{model}$ since calculating the expected value over the entire model would require knowing all possible configurations between visible and hidden states in order to obtain the probability of each one and compute the expected value [24]. This problem quickly becomes intractable [22] . For example, in an RBM with 100 nodes in the visible layer and 200 nodes in a hidden layer, there would be $2^{300}$ possible configurations, a number greater than the number of atoms in the universe ($10^{78}-10^{82}$ atoms in the universe). Storing all that information is impossible, even for a relatively simple network like the one described above.
@@
A much faster learning procedure called Contrastive Divergence was proposed by Hinton [23]. It starts by setting the states of the visible units to a training vector. Then, the binary states of the hidden units are all computed in parallel using Eq. 18. Once the binary states have been determined for the hidden units, a “reconstruction” is produced by setting each $v_{i}$ to 1 with a probability given by Eq. 19. The change in a weight is then given by:
@@
RBMs typically learn better models if more steps of alternating Gibbs sampling are used before collecting the statistics for the second term in the learning rule in Eq. 20 , which will be called the negative statistics [24].
@@
While Contrastive Divergence facilitates efficient learning in RBMs, their true power lies in their representational capacity, enabling them to capture a wide range of distributions. Increasing the number of hidden units enhances performance by improving training log-likelihood and reducing the KL divergence between data and model distributions [25]. Research has shown that an RBM with $k+$1 hidden units can closely approximate any distribution over $\{0,1\}^{n}$, where $k$ denotes the number of input vectors with non-zero probability. Furthermore, it has been demonstrated that an RBM with $12^{n}-$ 1 hidden units can approximate any distribution [26]. Refinements in these findings suggest that RBMs with $\alpha 2^{n}$-1 hidden units (where $\alpha<1$) are sufficient for approximating any distribution [27]. This versatility makes RBMs particularly promising for addressing challenges such as Full Configuration Interaction (FCI) in quantum mechanics, where the curse of dimensionality severely limits computational tractability for complex systems. By leveraging their ability to capture complex distributions efficiently, RBMs offer a potential solution to overcoming the computational challenges posed by FCI.

### v2

@@ B.   Background on FCI and RBM
To address this issue, our approach leverages an RBM to guide the sampling process. An RBM is an unsupervised generative model that learns complex, non-linear probability distributions over the configuration space. It consists of two layers of nodes: a visible layer that represents the observed data and a hidden layer that captures dependencies between variables (Figure 1).
Figure 1: General architecture of the Restricted Boltzmann Machine.
The model is defined by its energy function:
$E(v,h)=-\sum_{i}a_{i}v_{i}-\sum_{j}b_{j}h_{j}-\sum_{i,j}v_{i}h_{j}w_{ij},$ (4)
which determines a joint probability distribution over visible ($v$) and hidden ($h$) units:
$p(v,h)=\frac{1}{Z}e^{-E(v,h)},$ (5)
with $Z$ being the partition function (Zhang et al., 2018, Hinton, 2012). Here, $a_{i}$ and $b_{j}$ are the biases associated with visible and hidden units respectively, and $w_{ij}$ represents the weight connecting visible unit $i$ to hidden unit $j$.
RBMs are trained using Contrastive Divergence (Hinton, 2012), which enables efficient learning of the model parameters by approximating the gradient of the log-likelihood. This training procedure helps the RBM to quickly capture the underlying probability distribution of the input data without the need for exhaustive sampling of the partition function.
To sample from the model, the RBM alternates between updating hidden and visible units based on conditional activation probabilities. These are given by:
$p(h_{j}=1\mid v)$ $=\sigma\left(\beta\left(b_{j}+\sum_{i}v_{i}w_{ij}\right)\right),$ (6)
$p(v_{i}=1\mid h)$ $=\sigma\left(\beta\left(a_{i}+\sum_{j}h_{j}w_{ij}\right)\right),$ (7)
where $\sigma(x)=1/(1+e^{-x})$ is the logistic sigmoid function. These equations define the Gibbs sampling process used during training, where $\beta$ is the inverse temperature. This parameter controls the level of stochasticity: Low $\beta$ values (high temperatures) lead to more random behavior, while high $\beta$ values result in sharper, more deterministic activations.
The true strength of the RBM lies in its representational capacity. In theory, an RBM can learn to approximate a wide range of probability distributions. In fact, increasing the number of hidden units improves performance by enhancing the training log likelihood, and reducing the Kullback-Leibler divergence between the data distribution and the model’s distribution (Le Roux and Bengio, 2008). It has been shown that an RBM with $k+1$ hidden units can closely approximate any distribution over $\{0,1\}^{n}$, where $k$ denotes the number of input vectors with non zero probability. Moreover, it has been demonstrated that an RBM with $12^{n}-$ 1 hidden units can approximate any distribution (Montufar and Ay, 2011), and further refinements suggest that RBMs with $\alpha 2^{n}-1$ hidden units (with $\alpha<1$) are sufficient for a broad class of distributions (Montúfar and Rauh, 2017).
In the context of this work, such representational power is crucial: The RBM efficiently captures complex distributions over electronic configurations and enables targeted sampling of the most relevant determinants. By prioritizing configurations with higher contributions to the wave function, the RBM guided approach significantly reduces the computational cost of FCI calculations, compared to traditional Monte Carlo methods. For a more detailed discussion on the functioning and training of RBMs, please refer to Appendix A.
Previous works have utilized RBMs in quantum many-body problems, but with different objectives. Carleo et al. Carleo and Troyer (2017) employed RBMs to parameterize the wavefunction itself, approximating $\psi$ through the joint probability distribution of visible and hidden units (Equation 5). This approach does not generate new determinants, but instead represents the wavefunction compactly. Basile et al. Herzog et al. (2022) trained an RBM using a randomly sampled dataset from CISD to assist in determinant selection, incorporating explicit coefficient weights to prioritize determinants and removing highly dominant ones like Hartree-Fock. Both methods offer significant insights, yet face limitations: Carleo’s approach does not directly identify new determinants beyond the fixed basis, and Basile’s method relies heavily on biased coefficient sampling. Addressing these limitations motivates our proposed strategy, described next.
III.   Proposed Approach
This section details the implementation of our RBM-guided sampling strategy to optimize the sampling process in the FCI method. Building upon the RBM’s ability to model complex configuration distributions, we introduce an iterative sampling mechanism that progressively refines the determinant set. Unlike traditional methods such as Hartree–Fock or CISD, which rely on a fixed and often truncated determinant space, our method dynamically generates and expands the set of Slater determinants across virtual orbitals. By focusing on configurations with higher predicted contributions, each iteration yields increasingly accurate wavefunction approximations while avoiding redundant exploration.
For consistency in the initial setup and energy evaluation, we employ Quantum Package Garniron et al. (2019) to:
  1. 1.
    Compute the self-consistent Hartree–Fock wavefunction.
  2. 2.
    Obtain the CISD determinants, which serve as the initial ansatz for our iterative approach.
  3. 3.
    Perform diagonalization of the Hamiltonian matrix at each iteration using the generated determinants, employing the Davidson method to efficiently obtain the ground state eigenvalue.
The full algorithm is summarized in Algorithm 1. The process begins with an initial set of CISD determinants and iteratively refines the determinant pool through RBM-guided sampling. On each iteration, the RBM is trained to learn an approximate distribution over the configuration space. New determinants are then generated based on this learned distribution and combined with those from previous iterations.
After removing duplicates and previously discarded configurations, the Hamiltonian is built and diagonalized to obtain the CI coefficients. Determinants whose coefficients fall below a predefined threshold Pthreshold are pruned and stored in a taboo list to avoid re-selection in future iterations. The pruning of determinants is performed based on their squared absolute coefficient values ($|c_{i}|^{2}$), as these reflect the probability weight of each configuration in the wavefunction expansion. This physically motivated choice ensures that only the most relevant configurations, those significantly contributing to the probability distribution, are retained. Consequently, it avoids systematic overestimation of the ground state energy and ensures that the wavefunction remains accurate. Then the determinant pool is updated based on the remaining significant determinants, and the Hamiltonian matrix is re-evaluated with this new set of configurations. This process repeats until the total energy change between consecutive iterations falls below a stability threshold Sthreshold, indicating convergence. A high-level overview of the proposed approach is illustrated in Figure 2.
Figure 2: Overview of the RBM guided sampling and pruning cycle, including the use of a taboo list to avoid redundant configurations.
Each iteration refines the selection of relevant determinants, yielding an increasingly accurate approximation of the wavefunction. Importantly, as the configuration space grows with each iteration, the resulting wavefunction approaches the Full CI solution within the chosen basis. This property stems from the fact that the CI wavefunction spans a subspace of the Hilbert space, and its accuracy systematically improves as more determinants are included (Szabo and Ostlund, 1996b, David Sherrill and Schaefer, 1999, Shepherd et al., 2012).
Unlike other approaches that explicitly exclude highly dominant configurations from the training data, such as Hartree–Fock determinants to avoid over representation, our method does not impose any such bias. All determinants, regardless of their coefficient magnitude, are included during training. This allows the RBM to learn an unbiased probability distribution over the configuration space, leading to a more complete and physically representative model of the wavefunction.
As a result, our method achieves high quality approximations while using significantly fewer determinants than traditional approaches.
Algorithm 1 Generative Model with Determinant Pruning
  Define constants:
[…]
  $\text{Pthreshold}\in\mathbb{R}$Coefficient pruning threshold}
  $\text{Sthreshold}\in\mathbb{R}$Energy convergence threshold}
  $\text{epochs}\in\mathbb{N}$
  Initialize variables:
  $E\leftarrow 0$Initial energy}
  $\text{PDetsList}\leftarrow\emptyset$List of pruned determinants}
[…]
  for iteration $=1$ to max_iterations do
     for epoch $=1$ to epochs do
         Train RBM (dets)
     end for
[…]
     $\mathbf{D}^{\prime}\leftarrow\text{$\mathbf{D}^{\prime}$}+\text{dets}$
     $\mathbf{D}^{\prime}\leftarrow\text{Unique($\mathbf{D}^{\prime}$)}$
[…]
     $\mathbf{C}\leftarrow\text{Diagonalization($\mathbf{D}^{\prime}$)}$
[…]
     $E_{\text{new}}\leftarrow\text{ComputeEnergy}(\text{dets})$
     if $|E_{\text{new}}-E|<\text{Sthreshold}$ then
         break
     end if
     $E\leftarrow E_{\text{new}}$
  end for
A.   Innovations and Implementation Details
Our approach introduces several innovations to enhance the efficiency and relevance of determinant selection in the RBM-guided sampling process. First, we employ a taboo list to prevent the re-selection and re-diagonalization of previously pruned determinants, reducing redundant computations and significantly improving efficiency. Second, while the RBM learns the probability distribution of relevant determinants, some generated determinants may differ from those in the current set by more than two spin orbitals. According to Slater’s rules, the Hamiltonian matrix element for such determinants will be zero, meaning they do not directly connect to the existing determinant set, and therefore do not contribute meaningfully to the ground state energy. Our approach explicitly filters out these unconnected determinants by computing their Hamming distance with respect to the existing determinant set. Ensuring that only determinants with Hamming distances of at most four (allowing up to two differences per alpha and beta spin) are retained. This filtering step ensures a connected determinant space, allowing our model to reach an accurate solution with fewer determinants.
Furthermore, we go beyond determinant selection by analyzing the learned distributions across different molecules and basis sets, revealing patterns that were previously unexamined. To further optimize efficiency, we employ a binary-to-decimal conversion strategy Tubman et al. (2020) for determinant uniqueness verification, enabling fast vectorized operations and streamlined data handling.
Finally, to improve the selection process, we implement a non-uniform finite distribution sampler Krauth (2006) that biases electron permutations based on the RBM learned distribution. This ensures that determinant generation remains physically meaningful while optimizing computational performance.
By integrating these innovations, our method achieves high accuracy with significantly fewer determinants compared to traditional FCI or perturbatively selected CI methods like CIPSI. The next section evaluates the method’s performance, analyzing both its computational efficiency and the distribution it learns.
B.   Performance and Convergence Analysis
Our pruning method based on squared coefficients ($|c_{i}|^{2}$) systematically refines determinant selection, progressively enhancing the accuracy of the ground state energy approximation.
Furthermore, although previously pruned determinants are excluded from subsequent iterations via the taboo list, their future relevance cannot be completely dismissed. In principle, new determinants generated in later iterations might establish connections with previously discarded determinants, potentially increasing their importance in the wavefunction expansion. Nevertheless, this scenario is unlikely due to two main reasons: Firstly, we use a very low pruning threshold, ensuring that only configurations with negligible contributions are discarded; secondly, our iterative method naturally progresses towards generating determinants with higher order excitations, which typically have lower contributions (weights) to the overall wavefunction. Thus, any new connections involving previously pruned determinants would most likely involve high energy (low weight) configurations, making significant reconnection improbable.
As a result, our pruning strategy not only ensures accurate determinant selection at each iteration but also avoids unnecessary recomputation, contributing to overall computational efficiency. Rather than re-evaluating discarded configurations, each iteration focuses on refining the current selection, promoting a more effective exploration of the determinant space. This iterative refinement also keeps the computational cost of diagonalization within manageable limits, maintaining the expected $O(n^{2})$ or at worst $O(n^{3})$ scaling of the Davidson algorithm during the diagonalization step.
IV.   Results and Discussion
We evaluated our model on $H_{2}O$, $N_{2}$, and $C_{2}$ using the 6-31G and cc-pVDZ basis sets. Figure 3 shows the energy convergence across iterations, revealing that our method consistently approaches FCI energy levels. Notably, the final energies also fall within the range of CCSD(T) benchmarks, confirming that our model delivers chemically accurate results while being computationally more efficient.
Table 2 provides a detailed comparison between CISD, FCI, CCSD(T), and our approach. Across all systems, our method achieves an accuracy of at least 99.99% of the FCI energy after a small number of iterations, using significantly fewer determinants.
Molecule Basis CISD Energy FCI Energy Our Energy CCSD(T) Energy Accuracy w.r.t. FCI (%) Our Num Dets Iterations
H2O cc-pVDZ -76.232000 -76.243900 -76.243100 -76.243405 99.998 1304036 8
H2O 6-31G -76.113900 -76.122370 -76.120768 -76.120905 99.997 213662 8
C2 cc-pVDZ -75.666900 -75.729840 -75.726530 -75.727818 99.995 180068 9
C2 6-31G -75.584600 -75.644180 -75.642960 -75.792674 99.998 93400 8
N2 cc-pVDZ -109.243800 -109.280700 -109.268200 -109.276482 99.988 1728927 9
N2 6-31G -109.075900 -109.108420 -109.098980 -109.104365 99.991 806720 9
Table 2: Comparison of Energy Approximations and Computational Effort: A Comparative Analysis between CISD, FCI, CCSD(T), and Our Proposed Model, Including the Number of Determinants and Iterations Required for Energy Convergence.
For instance, in the case of C2 with the 6-31G basis, our model achieved 99.998% accuracy with only 93,000 determinants, while CIPSI required 938,000 to reach 99.999%. On average, our method reduced the number of determinants by 40.5% compared to CIPSI, with savings ranging from 30% to 50%.
We also compared our method with the results reported by Basile and Herzog Herzog et al. (2022), who required 19 million determinants to reach 99.979% of the correlation energy for N2 with the cc-pVDZ basis. Our model achieved a comparable accuracy of 99.98% using only 1.3 million determinants, representing a reduction by a factor of fifteen. Furthermore, under more conservative pruning settings, our approach reached the same 99.979% target using just 137,412 determinants, highlighting the efficiency and scalability of our training strategy and pruning mechanism.
We adjusted our model using a grid search, varying the number of epochs, temperature ($1/\beta$ in Eq. 22 and Eq. 23), batch size, and the number of Gibbs sampling steps used to approximate the hidden-visible distribution in the RBM. To illustrate the impact of temperature, Figure 4 shows the results for the $H_{2}O$ molecule using the cc-pVDZ basis as a representative example. Overall, the results indicate that the model performs best at a temperature of $11/\beta=$ 1, suggesting that this value allows the RBM to effectively learn the underlying distribution. In contrast, higher temperatures ($51/\beta=$ 5 and $301/\beta=$ 30), which introduce greater stochasticity, led to consistently worse performance. In particular, the setting $301/\beta=$ 30, which corresponds to random sampling, resulted in the largest deviation from the FCI energy.
Additionally, across all temperature settings, we observed that fewer epochs, smaller batch sizes, larger numbers of Gibbs sampling steps, and lower learning rates resulted in better performance compared to training for more epochs. This suggests that the model aims to approximate the overall distribution at each iteration, rather than overfitting to the current training set. Consequently, it maintains the ability to sample determinants beyond the training distribution, enhancing generalization. These findings are consistent with the patterns observed in Figure 4, which summarizes the performance of various hyperparameter combinations.
Figure 4: Hyperparameter search results (only 4 iterations) showing energy (x axis) versus epochs (y axis) for $H_{2}O$ molecule in the cc-pVDZ basis for different temperatures $51/\beta=1,$ 5 and $30$. Models with lower temperatures and fewer epochs performed best, approaching FCI energy more closely.
Analyzing the estimated probability distribution of each Molecular Orbital (MO) for the water molecule in the 6-31G basis, Figure 5 shows a transition from a Fermi-Dirac-like shape to a distribution predominantly concentrated in the lowest energy levels (first MOs for alpha and beta electrons), with a noticeable undulating pattern. This distribution is obtained by computing how frequently each MO is occupied during determinant generation at each iteration, and optionally weighting these frequencies by the squared coefficients ($|c_{i}|^{2}$) obtained from diagonalization. The resulting bar plots effectively capture the model’s learned sampling distribution for $H_{2}O$ in the 6-31G basis.
These results may also explain the poor performance of high temperature models in Figure 4. RBMs with completely random training moves have difficulty exploring multimodal distributions efficiently Béreux et al. (2023), leading to less effective sampling. This also clarifies why models trained with $11/\beta=$ 1 perform better with fewer epochs. Multimodal distributions are particularly challenging to sample because simulations can remain in metastable energy minima for extended periods, limiting the exploration of the configuration space. Training with fewer epochs helps mitigate this issue by reducing overfitting and improving coverage of relevant determinants.
Figure 5: Estimated probability distribution of orbital occupation (blue bars), and the same distribution weighted by the coefficients obtained from diagonalization (green bars) for $H_{2}O$ in the 6-31G basis at each iteration.
A similar trend appears when weighting the distributions by the coefficients obtained from diagonalization ($c_{i}$ in Eq. 1), shown in green in Figure 5. These coefficients correspond to the squared amplitudes $|c_{i}|^{2}$ of each determinant and can be interpreted as their physical contribution to the total wavefunction, thus offering a probability-based view of orbital occupation.
Interestingly, the RBM consistently learns similar orbital occupancy patterns across multiple runs, even when starting from random initialization. While differences appear across different molecules, the same molecule across different basis sets exhibits recurring structural patterns. This is further illustrated in Figure 6, where similar distributions emerge for $H_{2}O$ under different bases, while distinct molecules show more significant variations.
We verified that the RBM learns an underlying probability distribution by conducting two experiments on the same molecule. In one experiment, we randomly reinitialized the weights at each iteration, while in the other, the weights remained unchanged. Remarkably, both experiments produced distributions with the same behavior at each iteration (Figure 7). This suggests that the RBM is not merely memorizing configurations, but is instead consistently identifying probability distribution of electrons in Molecular Orbitals.
Figure 8 shows that both experiments exhibit very similar energy convergence towards the FCI energy, but the experiment without reinitialization shows slightly faster convergence at early iterations. This suggests that avoiding reinitialization may provide an advantage in convergence speed. This strong similarity further supports the idea that the observed undulating pattern is not a feedback artifact generated by the RBM, but rather a learned feature of the system itself.
Figure 7: Comparison between frequency occupation in each Molecular Orbital reinitialized and not reinitialized in each iteration for $H_{2}O$ in 6-31G basis for each iteration.
The data from Figures 6 and 7 indicate that the model consistently converges to the same distribution, regardless of the molecular basis or weight initialization. This behavior suggests ergodic sampling, as the distribution continuously adjusts at each iteration.
To formally verify that our sampling process is ergodic and does not produce non convergent Markov chains, we leverage the fact that the RBM is not trained via gradient descent, as in conventional neural networks, but instead through Gibbs sampling, a specific case of the Markov Chain Monte Carlo (MCMC) method Chib (2001). Given this formulation, we can apply the Gelman-Rubin statistic to assess convergence. Consider $j$ Markov chains, each initialized with different random values. For each chain $i$ of $L$ steps, the mean value of the chain is given by:
$\bar{x}_{j}=\frac{1}{L}\sum_{i=1}^{L}x_{i}^{(j)}$ (8)
The mean of the means across all chains is:
$\bar{\bar{x}}_{j}=\frac{1}{J}\sum_{j=1}^{J}\bar{x}_{j}$ (9)
The variance of the means of the chains, also known as the between chain variance $B$, is defined as:
$B=\frac{L}{J-1}\sum_{j=1}^{J}\left(\bar{x}_{j}-\bar{\bar{x}}_{j}\right)^{2}$ (10)
The average variance of the individual chains across all chains, known as the within chain variance $W$, is:
The Gelman-Rubin statistic $R$ is then defined as Vats and Knudson (2021):
$R=\sqrt{\frac{\frac{L-1}{L}W+\frac{B}{L}}{W}}$ (12)
In our case, since the different runs of the algorithm converge to similar distributions after a certain number of iterations, and the energy values are consistent across these runs, it is as if we were running $J$ simulations where the variance of the means of the chains is very low. For instance, the variance of the normalized distributions between the reinitialized and non-reinitialized weights is 3 ×$\times 10^{-4}$, which corresponds to the value of $B$ in the Gelman-Rubin equation. This low variance implies that $B$ is close to zero, leading to an estimate of the Gelman-Rubin statistic $R$ that is very close to 1. This result suggests that our sampling process is ergodic and not subject to non convergence issues.
Figure 8: Energy convergence for the experiments with reinitialized and not reinitialized weight on each iteration for $C_{2}$ in 6-31G basis.
Another analysis confirming the RBM’s ability to learn consistent distributions is shown in Figure 9. Gaussian functions were fitted to the estimated probability distributions from Figure 6 to highlight general patterns. The resulting curves display similar behavior, with their means centered around the same molecular orbitals and showing consistent shapes across different basis sets. Naturally, the cc-pVDZ basis spans a larger number of orbitals than 6-31G, which explains the broader range observed in the distribution.
Analyzing the behaviors of the envelope functions of Figure 9, a pattern emerges that resembles the radial probability distribution shown in Figure 10. This plot shows a radial probability density $\psi^{2}r^{2}$ against $r$, illustrating that electrons in higher energy orbitals are more likely to be found farther from the nucleus. The recurring modulations observed across different basis sets may suggest a connection to the molecule’s electronic configuration.
We also observe a similar pattern to the RDFs obtained from experiments, rigid models, and Monte Carlo simulations for different molecules. Specifically, in Figure 9, the learned distribution for water molecule ($H_{2}O$) exhibit two prominent peaks that closely match those observed in the RDF of water at 290K and 1 bar (Figure 11), generated using experimental and simulation data (McBride et al., 2012, Cordeiro, 1998). These peaks have been previously associated with the O-H bond structure. Similarly, for $C_{2}$, our model captures modulated patterns that align with RDF results reported in (Alavi and Thompson, 2004, Koperwas et al., 2020), while for $N_{2}$, the learned distributions show similarities with the RDF obtained under experimental conditions (Cassone et al., 2022). These observations reinforce the idea that the RBM effectively learns electron distribution patterns that reflect molecular bonding characteristics.
Figure 11: Radial Distribution Function (RDF) for water at 290K and 1 bar. The RDF was constructed using experimental and simulation data reported in (Cordeiro, 1998, McBride et al., 2012).
Finally, an important analysis is observed in Figure 12, where we compare the distribution learned by the RBM with the normalized orbital energies for each molecule in the cc-pVDZ basis. The orbital energies were computed using the software Q-chem (Shao, 2014). Remarkably, the phenomenon of having two or four prominent energy peaks in the lowest states is observed in the distribution learned by the models (Fig 9). This suggests that the model might be learning relevant information about orbital energy, particularly for those with the lowest energy levels.
Figure 12: Comparison of the frequency distribution learned by the RBM with the normalized orbital energies for each molecule in the cc-pVDZ basis. The RBM captures distinct energy peaks, particularly in the lowest energy states, showing a remarkable correspondence with the computed orbital energies using Q-Chem (Shao, 2014).
V.   Conclusion
We introduced an explainable and efficient method based on an RBM to approximate FCI calculations, significantly improving efficiency while preserving accuracy. Our approach represents an improvement over Quantum Monte Carlo (QMC) methods, offering a more structured and efficient way to explore determinant space. By effectively identifying and sampling the most relevant determinants, our model reduces the number of required determinants by up to four orders of magnitude compared to FCI and by 30–50% compared to CIPSI, without compromising precision. The obtained energy values consistently align with those from highly accurate quantum chemistry methods, such as CCSD(T), demonstrating the robustness of our approach.
Beyond its efficiency, our method provides a deeper understanding of the electronic structure. The RBM learns an intrinsic probability distribution over molecular determinants, capturing relevant quantum patterns. By analyzing the frequency of selected determinants, we observe a strong relationship between the learned distributions and molecular orbital occupancies, suggesting that the model is learning meaningful electronic structure features. Moreover, our results reveal undulating patterns in the learned determinant distributions that closely resemble Radial Distribution Functions (RDFs) associated with molecular bonding, further reinforcing the physical interpretability of our approach.
Additionally, the ergodic nature of our sampling process was validated through statistical analysis, confirming that our RBM-based method does not produce non convergent Markov chains. This ensures that the learned distributions remain consistent across multiple runs, further supporting the robustness of our methodology.
Our findings indicate that machine learning can play a crucial role in accelerating electronic structure calculations while preserving interpretability, offering a bridge between AI driven approaches and traditional quantum chemistry methods. Future work will explore extending this methodology to larger molecular systems and incorporating hybrid quantum and classical approaches to further improve efficiency and scalability.
CRediT authorship contribution statement
Sandra Leticia Juárez-Osorio: Review and Editing, Assistance in Software Development, and Support in Figure Preparation.
Acknowledgements
This work was partially supported by CONAHCYT, Mexico
Declaration of competing interest
Supplementary Information
A modular and fully reproducible implementation of the proposed method is available in Python. The source code and supplementary materials can be accessed at:
https://github.com/Navihdz/FCI-RBM
Appendix A: Detailed Explanation of Restricted Boltzmann Machines
@@ Acknowledgements
Restricted Boltzmann Machines is a type of generative model and unsupervised learning algorithm used in the field of machine learning. RBM is a probabilistic energy-based neural network consisting of two layers of nodes: a visible layer and a hidden layer. The nodes in the visible layer represent observed variables, while the nodes in the hidden layer represent latent or unobserved features. These two layers are fully connected to each other, but there are no connections within each layer (See Figure 13).
Figure 13: General architecture of the Restricted Boltzmann Machines.
@@
The main goal of an RBM is to learn the synaptic weights that maximize the joint probability of the observed data in the visible layer and the latent representations in the hidden layer, or in other words the RBM as a generative model, represents a probability distribution (over states of the visible units) where low energy configurations have higher probability. The energy is determined by the connection weights which are the parameters to be learnt from the data to built a generative model that learns the underlaying dynamic of the system (Zhang et al., 2018). This is achieved using an algorithm called “Gibbs sampling” or “Contrastive Divergence” (Hinton, 2002).
@@
Because there are no direct connections between hidden units in an RBM, it is not necessary to know all the system configurations but rather the probabilities that connect a visible state to each of the nodes in the hidden state. Therefore, obtaining a sample of $\langle v_{i}h_{j}\rangle_{data}$ is very easy. Given a randomly selected training input, $v$, the binary state, $h_{j}$, of each hidden unit, $j$, is set to 1 with the following probability (Hinton, 2012):
@@
The problem arises when trying to obtain the expectation under the model distribution $\langle v_{i}h_{j}\rangle_{model}$ since calculating the expected value over the entire model would require knowing all possible configurations between visible and hidden states in order to obtain the probability of each one and compute the expected value Hinton (2012). This problem quickly becomes intractable (Zhang et al., 2018) . For example, in an RBM with 100 nodes in the visible layer and 200 nodes in a hidden layer, there would be $2^{300}$ possible configurations, a number greater than the number of atoms in the universe ($10^{78}-10^{82}$ atoms in the universe). Storing all that information is impossible, even for a relatively simple network like the one described above.
@@
A much faster learning procedure called Contrastive Divergence was proposed by Hinton (Hinton, 2002). It starts by setting the states of the visible units to a training vector. Then, the binary states of the hidden units are all computed in parallel using Eq. 22. Once the binary states have been determined for the hidden units, a “reconstruction” is produced by setting each $v_{i}$ to 1 with a probability given by Eq. 23. The change in a weight is then given by:
@@
RBMs typically learn better models if more steps of alternating Gibbs sampling are used before collecting the statistics for the second term in the learning rule in Eq. 24 , which will be called the negative statistics (Hinton, 2012).
@@
While Contrastive Divergence facilitates efficient learning in RBMs, their true power lies in their representational capacity, enabling them to capture a wide range of distributions. Increasing the number of hidden units enhances performance by improving training log likelihood and reducing the KL divergence between data and model distributions (Le Roux and Bengio, 2008). Research has shown that an RBM with $k+$1 hidden units can closely approximate any distribution over $\{0,1\}^{n}$, where $k$ denotes the number of input vectors with non-zero probability. Furthermore, it has been demonstrated that an RBM with $12^{n}-$ 1 hidden units can approximate any distribution (Montufar and Ay, 2011). Refinements in these findings suggest that RBMs with $\alpha 2^{n}$-1 hidden units (where $\alpha<1$) are sufficient for approximating any distribution (Montúfar and Rauh, 2017). This versatility makes RBMs particularly promising for addressing challenges such as FCI in quantum mechanics, where the curse of dimensionality severely limits computational tractability for complex systems. By leveraging their ability to capture complex distributions efficiently, RBMs offer a potential solution to overcoming the computational challenges posed by FCI.